OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line?

OpenStudy (anonymous):

y − 4x = 1 y + 4x = 1 x − 4y = 1 x + 4y = 1

OpenStudy (anonymous):

@reneesweet17

OpenStudy (anonymous):

@SolomonZelman can yall help me please?

OpenStudy (anonymous):

im not good at math hun

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it negative or posative

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

i think it the only one with the Y then a plus sign

OpenStudy (anonymous):

do maybe B

OpenStudy (anonymous):

it was wrong smh

OpenStudy (anonymous):

sorry :( i tried